OpenStudy (anonymous):

Solve y = ln (1/e^3) huh???

jimthompson5910 (jim_thompson5910):

solve for x? or y?

OpenStudy (anonymous):

that's all the question said, these are the answer choices A. -3 B. -e C. e D. 3

jimthompson5910 (jim_thompson5910):

ok so solve for y

jimthompson5910 (jim_thompson5910):

hint: 1/e^3 = e^(-3)

jimthompson5910 (jim_thompson5910):

ln(x^y) = y*ln(x) ln(e) = 1

OpenStudy (anonymous):

so would y=-3

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

thanks!

jimthompson5910 (jim_thompson5910):

np